1. Suppose the population of Middletown is 56,000 and its population is increasing by 1/15 each year: A) How many people will reside in Middletown at the end of 5 years? B) How many years will it take for the population to reach 75,000 people?
Answer by stanbon(75887)

1. Suppose the population of Middletown is 56,000 and its population is increasing by 1/15 each year: A) How many people will reside in Middletown at the end of 5 years?
-------------------
A(5) = 56000(1+(1/15))^5
A(5) = 77327
-------------------------------
B) How many years will it take for the population to reach 75,000 people?
75000 = 56000(1+(1/15))^t
1.3393 = (1+(1/15))^t
---
Take the log and solve for "t":
t = 4.527 years
----------------------
Cheers,
Stan H.
==============